Lionel Corporation manufactures pharmaceutical products sold through a network of sales agents in the United States and Canada. The agents are currently paid an 18% commission on sales; that percentage was used when Lionel prepared the following budgeted income statement for the fiscal year ending June 30, 2019:

Lionel Corporation
Budgeted Income Statement
For the Year Ending June 30, 2019
($000 omitted)
Sales $ 28,700
Cost of goods sold
Variable $ 12,915
Fixed 3,444 16,359
Gross profit $ 12,341
Selling and administrative costs
Commissions $ 5,166
Fixed advertising cost 861
Fixed administrative cost 2,296 8,323
Operating income $ 4,018
Fixed interest cost 718
Income before income taxes $ 3,300
Income taxes (30%) 990
Net income $ 2,310

Since the completion of the income statement, Lionel has learned that its sales agents are requiring a 5% increase in their commission rate (to 23%) for the upcoming year. As a result, Lionel’s president has decided to investigate the possibility of hiring its own sales staff in place of the network of sales agents and has asked Alan Chen, Lionel’s controller, to gather information on the costs associated with this change.

Alan estimates that Lionel must hire eight salespeople to cover the current market area, at an average annual payroll cost for each employee of $80,000, including fringe benefits expense. Travel and entertainment expenses is expected to total $620,000 for the year, and the annual cost of hiring a sales manager and sales secretary will be $160,000. In addition to their salaries, the eight salespeople will each earn commissions at the rate of 10% of sales. The president believes that Lionel also should increase its advertising budget by $520,000 if the eight salespeople are hired.

Required

1. Determine Lionel’s breakeven point (operating profit = 0) in sales dollars for the fiscal year ending June 30, 2019, if the company hires its own sales force and increases its advertising costs. Prove this by constructing a contribution income statement.

2. If Lionel continues to sell through its network of sales agents and pays the higher commission rate, determine the estimated volume in sales dollars that would be required to generate the operating profit as projected in the budgeted income statement.
Answer #1

Present scenario:

Sales 28,700
Less: Variable costs
in COGS 12,915
in Commission (28700*0.18) 5166
Contribution 10,619
Less: Fixed Costs
in COGS 3,444
Advertisement 861
Administration 2296
Operating Profit 4,018
Interest (Fixed) 718
EBT 3,300
Taxes @ 30% 990
EAT (NET INCOME) 2310

a. Existing Fixed Costs : (from above) 861 + 2296 + 3444 + 718 =>7319

Additional Fixed costs under proposed method:

Payroll cost of staff 80,000 * 8 640000
Travel and entertainment Given 620000
Cost of sales manager and secretary Given 160000
Advertising 520000
Total 1940000
Omitting '000 1940

So, total Fixed Costs = 7319 + 1940 = 9259

Contribution under proposed method:

Sales (given) 28700
Variabe costs in COGS 12915
Commission (10%) 2870
Contribution is $ 12915
Contribution in % 12915/28700*100 = 45%

Hence BEP (Sales) = Fixed Costs/Contribution (%) => 9259/0.45 = $ 20575.56

Income statement(contribution approach):

BEP Sales 20575.56
LESS: variable costs
in COGS (20575.56*45%) 9259.00
Commission (10%) 2057.56
Contribution 9259.00
Fixed Costs (as computed above) 9259.00
Operating Profit 0.00

b, As per budget income statement, operating profit before tax after fixed interest payment is $ 3300. Total Fixed costs (existing) as in the beginning is $ 7319. Hence the required contribution is $ 10619. (ie., 3300+7319).

Now Contribution as per higher commission will be:

Sales 28,700
Less: Variable costs
in COGS 12,915
in Commission (28700*0.23) 6601
Contribution in Amount 9,184
Contribution in Percentage 32%

So Required Sales for same operating profit = Contribution reqd / Contribution % => 10619/0.32 = $ 33184.38 .